Find the equation of a circle which passes through the origin and whose center
is at the point (-1,1).

Answer by ryanman(19)

(x-h)^2 + (y-k)^2 = r^2
(x+1)^2 + (y-1)^2 = 2
by drawing the circle and using the pythagroem theroem , the radius = sqrt 2